At the end of April, Cavy Company had completed Job 766 and 765. According to the individual job cost sheets the information is as follows:

Job

Direct Materials

Direct Labor

Machine Hours

    

Job 765

$5,670

$3,500

27

Job 766

$8,900

$4,775

44

    

Job 765 produced 152 units, and Job 766 consisted of 250 units.

Assuming that the predetermined overhead rate is applied by using machine hours at a rate of $200 per hour, determine the (a) balance on the job cost sheets for each job, and (b) the cost per unit at the end of April.
Answer #1

a) Balance of Job cost sheet

Job 765 Job 766
Direct material 5670 8900
Direct labor 3500 4775
Overhead 27*200 = 5400 44*200 = 8800
Total Job cost 14570 22475

b) Cost per unit

Job 765 = 14570/152 = 95.86

Job 766 = 22475/250 = 89.90